OpenStudy (anonymous):

Which expression is equivalent to (7^3)-2? 1. 1 ------------- 7*7*7*7*7*7 2. 7 3. 1 ------------------- 6*6*6*6*6*6*6 4. 1 __ 7

OpenStudy (anonymous):

@Miyuru @SheldonEinstein @hero @ivano_logy

OpenStudy (anonymous):

@Miyuru

OpenStudy (anonymous):

\(\large (7^3)^{-2}=\frac{1}{(7^3)^2}= \) ?????